OpenStudy (anonymous):

what is the value of f(-2) f(x)= -2 if x<-3 4x if -3<(equal to) x <( equal to) -1 x^2 if x>-1 a.-8 b.4 c-2 d8

OpenStudy (cathyangs):

So for f(-2), x = -2. Which of those criteria (if x <-3, -3<_x<_-1, and -1<x) does x=-2 fall under in this piecewise function?

OpenStudy (anonymous):

the first one?

OpenStudy (anonymous):

-2 is x <-3?

OpenStudy (cathyangs):

No, -2 is actually -3<_x<_-1. When x = -2, it is greater than -3 and less than -1. Do you understand this?

OpenStudy (anonymous):

yes now i do thank you

OpenStudy (retireed):

it always helps to graph the piece wise equations. What is f(x) or y when x is less than -3 ? y = -2 for every x less than 2 What is f(x) or y when x is equal to or between -3 and -1 ? y = 4x for the stated domain What is f(x) or y when x is greater than -1 ? y = x^2 for every x greater than -1

OpenStudy (retireed):

x f(x) domain -10 -2 f(x)=-2 for x<-3 -9 -2 f(x)=-2 for x<-3 -8 -2 f(x)=-2 for x<-3 -7 -2 f(x)=-2 for x<-3 -6 -2 f(x)=-2 for x<-3 -5 -2 f(x)=-2 for x<-3 -4 -2 f(x)=-2 for x<-3 -3 -12 f(x)=4x for -3 <= x <= -1 -2 -8 f(x)=4x for -3 <= x <= -1 -1 -4 f(x)=4x for -3 <= x <= -1 0 0 f(x)=x^2 for x > -1 1 1 f(x)=x^2 for x > -1 2 4 f(x)=x^2 for x > -1 3 9 f(x)=x^2 for x > -1 4 16 f(x)=x^2 for x > -1 5 25 f(x)=x^2 for x > -1 6 36 f(x)=x^2 for x > -1 I think it would look like this (roughly)